find the vertix of the parabola
f(x)=25-10x+20x^2

Answer by Alan3354(69443)

Find the Line of Symmetry, x = -b/2a
x = 10/40
x = 1/4
---------
f(1/4) = 25 - 10/4 + 20/16
f(1/4) = 95/4
-----------
Vertex at (1/4,95/4)
or (0.25,23.75)